OpenStudy (anonymous):

What is the solution to the following system of equations? y = x2 + 5x + 2 y = x2 + 3x – 2 (–2, –4) (5, 3) (–2, 5) (3, 2)

OpenStudy (anonymous):

Solve like: x^2+5x+2=x^2+3x-2 and the substitute the value of x you obtain in any of the equations to get y. Then , you have to solve 5x+2=3x-2-----> x=-2 y=(-2)^2+5(-2)+2=4-10+2=4-8=-4 Try with the other equation and check it gives the same value for y when x=-2